# Interference-induced suppression of particle emission from a Bose-Einstein condensate in lattice with time-periodic modulations

### v1

@@ IV Summary
In a previous work where the configuration was somewhat similar, the collective modes resulted in a dramatic enhancement such that a small drive could induce large pulses when the drive frequency was tuned to resonance [33]. In the present model, however, the couplings give rise to a broader band, with only few particles ejected even when the drive strengths are increased, i.e., the emission is greatly suppressed. Since the central sites are coupled with both leads, this phenomenon might be attributed to the interference of the outcoming matter waves from the central sites, where the decay rate of the particles is much lower than the growth rate, leading to insignificant particle jets outside of the trap.
The geometries and the couplings can be readily implemented with optical lattices in experiments. One could, based on particular purposes, utilize specific engineering to either enhance or suppress the emission processes, and such framework would enable an intensive study of the correlations among the particle emission, the couplings and the various leads.
@@ Acknowledgements
L.Q.L would like to thank Professor Erich J. Mueller for the profound instructions. This work was supported by the Natural Science Research Start-up Foundation of Recruiting Talents of Nanjing University of Posts and Telecommunications (Grant No. NY223065), the China Scholarship Council (Grant No. 201906130092), and Natural Science Foundation of Sichuan Province (Grant No. 2023NSFSC1330).
@@ Appendix A Derivation of the analytical relation
$\partial_{t}v$ $=$ $*\displaystyle-\frac{g}{2}\left[e^{i\omega t}(2|u|^{2}+|v|^{2})v+e^{i(\omega+2\nu_{a}-2\nu_{s})t}u^{2}v^{*}\right.$ (53)
@@
$\partial_{t}|u|^{2}$ $=$ $\Gamma|u|^{2}-\frac{g}{2}\left[e^{i(\omega+2\nu_{s}-2\nu_{a})t}v^{2}u^{*2}\right.$ (54)
@@
$\left.-e^{i(\omega+2\nu_{s}-2\nu_{a})t}u^{*2}v^{2}-e^{-i(\omega+2\nu_{s}-2\nu_{a})t}u^{2}v^{*2}\right],$
with $\Gamma=4t_{c}^{2}{\rm{Im}}G_{11}(\nu_{s})$. A nice simplifying assumption is to treat $u$ and $v$ as real, we find
@@
where $X(t)=\sin(\omega t)\sin(4t_{ab}t)$. The above equations can readily be numerically solved. Conversely, one could also rewritten from Eqs. (43) and (44) that

### v2

@@ IV Summary
In a previous work where the configuration was somewhat similar, the collective modes resulted in a dramatic enhancement such that a small drive could induce large pulses when the drive frequency was tuned to resonance [31]. In the present model, however, a much broader band emerges with only few particles ejected even when the drive strengths are increased, i.e., the emission is greatly suppressed. This phenomenon might be attributed to the destructive interference of the outcoming matter waves from the central sites, leading to the decay rate of the particles being much lower than the growth rate, and there are insignificant particle jets outside of the trap.
The symmetric geometry and the couplings can be readily implemented with optical lattices in experiments. One could, based on particular purposes, utilize specific engineering to either enhance or suppress the dynamic processes, and such framework would enable an intensive study of the correlations among the particle emission, the couplings and the various leads. The ability of tailoring the dynamical excitation and coherently steering particles to produce directed motion in a lattice system also points to the possible implementation in coupled quantum dots and quantum information processing.
@@ Acknowledgements
L.Q.L. would like to thank Professor Erich J. Mueller for the profound instructions. This work was supported by the China Scholarship Council (Grant No. 201906130092), the Natural Science Research Start-up Foundation of Recruiting Talents of Nanjing University of Posts and Telecommunications (Grant No. NY223065), and the Natural Science Foundation of Sichuan Province (Grant No. 2023NSFSC1330).
@@ Appendix A Derivation of the analytical relation
$\partial_{t}v$ $=$ $-\frac{g}{2}\left[e^{i\omega t}(2|u|^{2}+|v|^{2})v+e^{i(\omega+2\nu_{a}-2\nu_{s})t}u^{2}v^{*}\right.$ (53)
@@
$\partial_{t}|u|^{2}$ $=$ $4t_{c}^{2}{\rm{Im}}G_{11}(\nu_{s})|u|^{2}-\frac{g}{2}\left[e^{i(\omega+2\nu_{s}-2\nu_{a})t}v^{2}u^{*2}\right.$ (54)
@@
$\left.-e^{i(\omega+2\nu_{s}-2\nu_{a})t}u^{*2}v^{2}-e^{-i(\omega+2\nu_{s}-2\nu_{a})t}u^{2}v^{*2}\right].$
A nice simplifying assumption is to treat $u$ and $v$ as real, we find
@@
where $X(t)=\sin(\omega t)\sin(4t_{ab}t)$. The above equations can readily be numerically solved. Conversely, one could also rewrite from Eqs. (43) and (44) that